OpenStudy (anonymous):

Two birds sit at the top of two different trees. The distance between the first bird and a bird-watcher on the ground is 14.1 feet. The distance between the bird-watcher and the second bird is 27.9 feet. What is the angle measure, or angle of depression, between this bird and the bird-watcher?

OpenStudy (anonymous):

30.4 59.6 63.2 87.9

OpenStudy (love_ranaa):

30.4

OpenStudy (bibby):

sin(x)=opposite/hypotenuse sin(x)=14.1/27.9 \(sin^{-1}(sin(x))=sin^{-1}(\dfrac{14.1}{27.9})\)

OpenStudy (bibby):

and so x = arcsin (or sin^-1) of 14.1/27.9

OpenStudy (anonymous):

so 30.4?

OpenStudy (bibby):

lol, just use a calculator http://www.wolframalpha.com/input/?i=arcsin(14.1%2F27.9)&t=crmtb01

OpenStudy (anonymous):

i got 59.6

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

I got 30.36

OpenStudy (bibby):

it's right there on the link I posted http://puu.sh/gnTOG.png

OpenStudy (anonymous):

okay thank you for help. i am really bad at using he calculators